OpenStudy (anonymous):

solve for x: (3-4x) over (5x) = -(5) over (6)

jhonyy9 (jhonyy9):

3-4x ----- = -5/6 5x 6(3-4x)= -5(5x) 18-24x= -25x 25x -24x = -18 x = -18